# Moments of inertia of rare-earth nuclei and the nuclear time-odd mean fields within exact solutions of the adiabatic theory

### v2

@@ Abstract
We systematically analyse the nuclear moments of inertia determined within the Skyrme and Gogny density functional theories. The time-odd mean fields generated by collective rotation are self-consistently determined by a novel exact iterative solution of the adiabatic time-dependent Hartree-Fock-Bogoliubov (ATDHFB) equations. Although details of the results depend on the functional used, the calculated moments of inertia are in good overall agreement with the experimental data, with no adjustable parameters. To show the essential importance of the time-odd mean fields, we compared the ATDHFB moments of inertia with those obtained from the Inglis-Belyaev formula. For Skyrme density functionals, we find strong correlations between the effective mass and the impact of the time-odd mean fields on the inertia’s rotational and vibrational moments.
@@
moment of inertia , density functional theory , adiabatic approximation , effective mass
@@
An exact determination of the ATDHFB inertia requires an inversion of the two-body stability matrix [2]. For deformed and/or superfluid systems, such explicit inversion requires a prohibitive computational cost because of the large matrix dimensions involved. Due to such numerical difficulties, the ATDHFB method was often used in the cranking approximation [3], which neglects the dynamical residual interactions generated by the collective motion. Collective inertia obtained in the cranking approximation is equivalent to that calculated using the Inglis [4] or Inglis-Belyaev (IB) [5] formula. The situation is further worsened in the perturbative cranking approximation [6, 3], where the inertia is calculated via the energy-weight moment tensors involving only the diagonal element of the stability matrix. Such an omission of the residual interaction, especially in the time-odd mean-field channels, leads to the violation of the Galilean symmetry that has severe consequences, for example, giving the wrong translational mass, see, e.g., Ref. [7].
@@
Despite the inconsistency caused by neglecting the time-odd mean fields, the cranking approximation was used in the theoretical studies of the collective inertia, from the calculation of the low-energy spectra [8] to the evaluation of nuclear fission half-life [9]. Such an inconsistency led to the enhancement factor adopted to correct the moment of inertia obtained from the cranking approximation [10]. To alleviate the inconsistency, a method based on the expansion of the inertia matrix was developed to solve the ATDHFB equation in the absence of pairing, which can evaluate the rotational moment of inertia with high accuracy of the order of 1% [11]. Recently, the similarity of the ATDHFB method and the finite-amplitude method at zero frequency was realised, and a quasiparticle finite-amplitude method on top of the Skyrme energy density functional was developed to evaluate the Thouless-Valatin rotational moment of inertia [12]. With the pairing strength adjusted to the experimental pairing gaps, the rotational moment of inertia in several axially deformed nuclei from the rare-earth and heavy-actinide regions was well reproduced, and the large and mass-dependent deviations from the values evaluated by the cranking approximation was confirmed. The finite-amplitude method was also used in Ref. [13] for the full five-dimensional quadrupole adiabatic inertia.
@@
In the present Letter, we developed a novel iterative method to solve the ATDHFB equations exactly and efficiently. We implemented the method within the non-relativistic DFT and determined the adiabatic time-odd mean-fields self-consistently. Below, we present a systematic analysis of the nuclear moments of inertia and a preliminary study of vibrational mass parameters in heavy deformed nuclei.
@@
The iterative ATDHFB method is sketched as follows [^1] We presented details of the iterative ATDHF method (with no pairing) in the recent conference publication [14]. In the nuclear DFT with pairing, the many-body wavefunction of the nuclear system is mapped onto a one-body quasiparticle density $\mathcal{R}$, whose time-evolution obeys the time-dependent Hartree-Fock-Bogoliubov equation, [2, 15],
$i\dot{\mathcal{R}}=[\mathcal{H},\mathcal{R}].$ (1)
@@
The one-body quasiparticle Hamiltonian $\mathcal{H}$ is the derivative of the energy of the nuclear system over the density. In the non-relativistic DFT, it is usually determined using either the zero-range Skyrme  [16] or finite-range Gogny [17] functional.
In the adiabatic limit, the velocity of the collective motion is assumed to be small enough such that the nucleus is almost in static equilibrium at all times. In such a case, the time-dependent equation (1) reduces to the ATDHFB equation,
$i\dot{\mathcal{R}}_{0}=i\dot{q}\frac{\partial\mathcal{R}}{\partial{q}}=[\mathcal{H}_{0},\mathcal{R}_{1}]+[\mathcal{H}_{1},\mathcal{R}_{0}],$ (2)
where $\mathcal{R}_{0}(q)$ and $\mathcal{H}_{0}(q)$ are, respectively, the static (time-even) density and Hamiltonian along the collective path parameterised by the collective variable $q$. The corresponding time-odd density $\mathcal{R}_{1}$ and Hamiltonian $\mathcal{H}_{1}$ are small corrections generated by the collective motion, and $\dot{q}$ denotes the collective velocity. For the collective rotation, the velocity is proportional to the angular momentum operator [18]; for the collective vibration, it can be evaluated from numerical differentiation [14]. Since up to the first order $\mathcal{H}_{1}$ depends linearly on $\mathcal{R}_{1}$, the ATDHFB equation (2) is a linear equation for $\mathcal{R}_{1}$ of many dimensions, which requires an unmanageable inversion of the huge stability matrix [2].
In the novel method proposed here to solve the ATDHFB equation (2), we transform the equation into an iterative fixed-point equation,
$i\dot{q}\frac{\partial\mathcal{R}}{\partial{q}}=[\mathcal{H}_{0},\mathcal{R}_{1}^{(n)}]+[\mathcal{H}_{1}^{(n-1)},\mathcal{R}_{0}],$ (3)
@@
where $\mathcal{H}_{1}^{(n-1)}$ is evaluated for $\mathcal{R}_{1}^{(n-1)}$ and $\mathcal{R}_{1}^{(0)}\equiv\mathcal{H}_{1}^{(0)}\equiv 0$. Then, at each iteration $n=1,2,\ldots{N}$, the collective inertia can be evaluated as
@@
$\mathcal{M}^{(n)}=\frac{i}{2\dot{q}}\textrm{Tr}(\frac{\partial\mathcal{R}}{\partial{q}}[\mathcal{R}_{0},\mathcal{R}_{1}^{(n)}]),$ (4)
@@
where the antisymmetric matrices $Z=\varphi^{\dagger}\mathcal{R}_{1}\chi$, $F=\varphi^{\dagger}\frac{\partial\mathcal{R}_{0}}{\partial q}\chi$, and $E_{1}=\varphi^{\dagger}\mathcal{H}_{1}\chi$ are defined by the standard quasiparticle and quasihole wave functions $\chi$ and $\varphi$, respectively, having the form
$\chi=\bigg{(}\begin{array}[]{c}A\\ B\\ \end{array}\bigg{)},\quad\varphi=\bigg{(}\begin{array}[]{c}B^{*}\\ A^{*}\\ \end{array}\bigg{)}.$ (6)
@@
Iteration of Eq. (5) can now proceed in a close analogy to the standard iteration of the HFB equation. Indeed, first we perform the SVD decomposition of the matrix $Z=U\Omega V^{\dagger}$, where $U$ and $V$ are unitary and $\Omega$ is diagonal and non-negative, i.e., $\Omega_{\mu\nu}=\delta_{\mu\nu}\omega_{\mu}$ and $\omega_{\mu}\geq 0$. Next, we find that the eigenequations for array $\mathcal{Y}=\left(\begin{array}[]{cc}0&Z\\ Z^{\dagger}&0\end{array}\right)$ (the first-order density matrix $\mathcal{R}_{1}$ expressed in the quasiparticle basis) read
$\mathcal{Y}\bigg{(}\begin{array}[]{c}U\\ \pm V\end{array}\bigg{)}=\bigg{(}\begin{array}[]{c}\pm U\Omega\\ V\Omega\\ \end{array}\bigg{)}=\pm\bigg{(}\begin{array}[]{c}U\\ \pm V\\ \end{array}\bigg{)}\Omega,$ (7)
that is, the eigenstates of $\mathcal{Y}$ appear in pairs of opposite eigenvalues. This compares perfectly well with the standard eigenstates of $\mathcal{R}_{0}$ (6), which also appear in pairs with eigenvalues 1 and 0. Therefore, with a minor change of adding factors $\omega_{\mu}$ to the summations of the quasiparticle wave functions, the standard HFB iteration can be ported to the ATDHFB iteration. Details of the procedure will be presented in the forthcoming publication [18].
@@
We see that the iterative equation (3) involves only one-body quasiparticle operators and the stability matrix does not need to be evaluated or inverted. Obviously, in the first iteration, in which $\mathcal{H}_{1}$ is neglected, $\mathcal{M}^{(1)}$ gives the IB moment of inertia. The method was implemented in the non-relativistic density functional solver HFODD [20, 21], which solves the nuclear DFT equations by expanding the quasiparticle wavefunction on a three-dimensional Cartesian harmonic-oscillator basis and allows for the self-consistent description of nuclei with arbitrary shapes.
The IB collective inertia has been widely used in multiple studies of nuclear collective motion. However, very early it was realized that the calculated values were too low to describe data. As a result, in most applications, the IB values were multiplied by ad hoc factors of about 1.3, see, e.g., recent Refs. [22, 23, 24] and references cited therein. Those were purely empirical factors meant to compensate for the unknown effects of the time-odd mean fields. In the present work, having a rapid and efficient method to include those effects, we microscopically determined the ratios of the ATDHFB and IB inertia and compared the full ATDHFB results with the data.
To this end, we performed systematic ATDHFB calculations for the Gd, Dy, Er, Yb, Hf, W, and Os isotopes with neutron numbers between 82 to 126. Two different Skyrme functionals, SkM* [25] and UNEDF1 [26], and one Gogny functional D1S [27] were used for this purpose. For SkM*, the volume pairing with the neutron and proton strengths of --178.83 and --211.20 MeV fm3 [12] were used, respectively. For UNEDF1, the mixed pairing of the type $V_{0}^{t}(1-\rho/\rho_{\textrm{sat}})$ was used with the neutron and proton parameters $V_{0}^{n}=-223.278$ and $V_{0}^{p}=-247.896$ MeV fm3, and with $\rho_{\textrm{sat}}=0.32~{}\textrm{fm}^{-3}$. Since the time-odd sectors of those Skyrme functionals were originally not adjusted to data, here we fixed them by employing the adjustments performed in Ref. [28].
@@
In Fig. 2, the distribution of the residuals between the calculated moments of inertia and the experimental values $(\mathcal{I}^{\textrm{Th.}}-\mathcal{I}^{\textrm{Expt.}})$ of all the experimentally available isotopes are presented for SkM*, UNEDF1, and D1S functionals, respectively. The shift between the distribution of the residuals for IB and ATDHFB moments of inertia can be clearly observed, which shows the importance of the time-odd mean field. Sudden increase of the deformation in the Hf, W and Os isotopes around $N=100$, predicted by UNEDF1 and D1S, is not visible in the experimental data. Therefore, these nuclei are excluded from the statistical analysis of the residuals between theoretical and experimental values. For SkM*, the mean value of the residual between the IB moment of inertia and the experimental data is -2.58 2superscriptPlanck-constant-over-$\hbar^{2}$/MeV, that is, for this functional, the IB formula underestimates the moment of inertia. On the contrary, an overestimation is predicted by the ATDHFB method with a mean value of 1.89 2superscriptPlanck-constant-over-$\hbar^{2}$/MeV. In both cases, the distributions are fairly narrow with standard deviations of 4.06 2superscriptPlanck-constant-over-$\hbar^{2}$/MeV (IB) and 3.62 2superscriptPlanck-constant-over-$\hbar^{2}$/MeV (ATDHFB). For D1S, the IB formula also underestimates the moment of inertia while the ATDHFB method overestimates them, but the distributions of the residual are wider than for SkM*. For UNEDF1, wider distributions are also found but both the IB formula and ATDHFB method overestimate the experimental values.

### v3

@@ Abstract
We systematically analyse the nuclear moments of inertia determined within the Skyrme and Gogny density functional theories. The time-odd mean fields generated by collective rotation are self-consistently determined by a novel exact iterative solution of the adiabatic time-dependent Hartree-Fock-Bogoliubov (ATDHFB) equations. Although details of the results depend on the functional used, the calculated moments of inertia are in good overall agreement with the experimental data, with no adjustable parameters. To show the essential importance of the time-odd mean fields, we compared the ATDHFB moments of inertia with those obtained from the Inglis-Belyaev formula. For Skyrme density functionals, we find strong correlations between the effective mass and the impact of the time-odd mean fields on the rotational and vibrational collective inertia.
@@
rotational moment of inertia , vibrational collective inertia , density functional theory , adiabatic approximation , effective mass
@@
An exact determination of the ATDHFB inertia requires an inversion of the two-body stability matrix [2]. For deformed and/or superfluid systems, such explicit inversion requires a prohibitive computational cost because of the large matrix dimensions involved. Due to such numerical difficulties, the ATDHFB method was often used in the cranking approximation [3], which neglects the dynamical residual interactions generated by the collective motion. Collective inertia obtained in the cranking approximation is equivalent to that calculated using the Inglis [4] or Inglis-Belyaev (IB) [5] formula. The situation is further worsened in the perturbative cranking approximation [6, 3], where the inertia is calculated via energy-weighted moment tensors that involve only the diagonal element of the stability matrix. Such an omission of the residual interaction, especially in the time-odd mean-field channels, leads to a violation of the Galilean symmetry, which has severe consequences, for example, resulting in the incorrect translational mass, see, e.g., Ref. [7].
@@
Despite the inconsistency caused by neglecting the time-odd mean fields, the cranking approximation has been used in theoretical studies of collective inertia, ranging from calculations of low-energy spectra [8] to evaluations of nuclear fission half-lives [9]. Such an inconsistency led to the enhancement factor adopted to correct the collective inertia obtained from the cranking approximation [10]. To alleviate the inconsistency, a method based on the expansion of the inertia matrix was developed to solve the ATDHFB equation in the absence of pairing, which can evaluate the rotational moment of inertia with high precision of the order of 1% [11]. Recently, the similarity of the ATDHFB method and the finite-amplitude method (FAM) at zero frequency was realised, and a quasiparticle FAM on top of the Skyrme energy density functional was developed to evaluate the Thouless-Valatin rotational moment of inertia [12]. With the pairing strength adjusted to the experimental pairing gaps, the rotational moment of inertia in several axially deformed nuclei from the rare-earth and heavy-actinide regions was well reproduced, and the large and mass-dependent deviations from the values evaluated by the cranking approximation were confirmed. The FAM was also used in Ref. [13] for the full five-dimensional quadrupole adiabatic inertia.
@@
This letter presents a novel iterative method for solving the ATDHFB equations exactly and efficiently. We implemented this method within the non-relativistic nuclear DFT framework and determined the adiabatic time-odd mean fields self-consistently. Below, we present a systematic analysis of the nuclear moments of inertia and a preliminary study of vibrational inertia in heavy deformed nuclei.
@@
The iterative ATDHFB method is outlined as follows. [^1] We presented details of the iterative ATDHF method (without pairing) in our recent conference publication [(Sun25a)]. In the nuclear DFT with pairing, the many-body wavefunction of the nuclear system is mapped onto a one-body quasiparticle density $\mathcal{R}$, whose time-evolution obeys the time-dependent Hartree-Fock-Bogoliubov equation [2, 15],
$i\hbar\dot{\mathcal{R}}(t)=\left[\mathcal{H}(t),\mathcal{R}(t)\right],$ (1)
where the dot represents the derivative over time, ˙$\dot{\mathcal{R}}(t)={\mbox{d}\mathcal{R}(t)}/{\mbox{d}{}t}$. We see that the static (time-independent) solution $\mathcal{R}_{0}$ then implies $\left[\mathcal{H}_{0},\mathcal{R}_{0}\right]=0$.
@@
At every time $t$, the one-body quasiparticle Hamiltonian $\mathcal{H}(t)=\partial{}E[\mathcal{R}(t)]/\partial\mathcal{R}(t)$ is the derivative of the energy $E[\mathcal{R}(t)]$ of the nuclear system over the generalised density matrix $\mathcal{R}(t)$. In non-relativistic nuclear DFT, it is typically determined using either the zero-range Skyrme functional [16] or the finite-range Gogny functional [17].
In the adiabatic limit, we assume that the system progresses in time along the given collective path $q(t)$, namely $E[\mathcal{R}(t)]\equiv{}E[\mathcal{R}(q(t))]$. Here, for simplicity, we consider only one collective coordinate $q$. In addition, we assume that the velocity $\dot{q}(t)$ of the collective motion is so small that the nucleus remains almost in static equilibrium at all times, that is,
$\mathcal{R}(q(t))\simeq\mathcal{R}_{0}(q(t))+\dot{q}(t)\tilde{\mathcal{R}}_{1}(q(t))+\ldots,$ (2)
$\mathcal{H}(q(t))\simeq\mathcal{H}_{0}(q(t))+\dot{q}(t)\tilde{\mathcal{H}}_{1}(q(t))+\ldots,$ (3)
$\left[\mathcal{H}_{0}(q(t)),\mathcal{R}_{0}(q(t))\right]\simeq 0,$ (4)
where the equalities hold only up to terms linear in collective velocity $\dot{q}(t)$. All these densities and mean fields depend on time only parametrically; that is, through the dependence on time of $q(t)$. This allows us to omit, for clarity, the implicit arguments $q(t)$ from the expressions that follow. In particular, the ATDHFB equation, which is the first-order part of the TDHFB equation (1), takes the following form,
$i\hbar\frac{\partial\mathcal{R}_{0}}{\partial{q}}=[\mathcal{H}_{0},\tilde{\mathcal{R}}_{1}]+[\tilde{\mathcal{H}}_{1},\mathcal{R}_{0}],$ (5)
where the derivative over $q$ replaced the time derivative, $\dot{\mathcal{R}_{0}}=\dot{q}\frac{\partial\mathcal{R}_{0}}{\partial{q}}$, and thus the ATDHFB equation became independent of the velocity $\dot{q}$.
Within the adiabatic approximation [1], the quasi-static zero-order densities and mean field are assumed to be time-even, $\mathcal{T}^{+}\mathcal{R}_{0}\mathcal{T}=\mathcal{R}_{0}$ and $\mathcal{T}^{+}\mathcal{H}_{0}\mathcal{T}=\mathcal{H}_{0}$. In contrast, the first-order corrections are assumed to be time-odd, $\mathcal{T}^{+}\tilde{\mathcal{R}}_{1}\mathcal{T}=-\tilde{\mathcal{R}}_{1}$ and $\mathcal{T}^{+}\tilde{\mathcal{H}}_{1}\mathcal{T}=-\tilde{\mathcal{H}}_{1}$.
For the collective rotation, the collective path is straightforward. It corresponds to the deformed state rotated in space by angle $\theta$, that is, $|\Phi(\theta)\rangle=\exp\left(i\theta\hat{I}_{x}\right)|\Phi(0)\rangle$, about the selected axis $x$, and then using $q\equiv\theta$. In this case, the density derivative $\frac{\partial\mathcal{R}_{0}}{\partial{q}}$ is given by the angular momentum $\hat{I}_{x}$ matrix elements, which can be calculated analytically [18]. For the collective vibration, it can be evaluated by the numerical differentiation of the HFB densities [(Sun25a)] obtained by constraining, e.g., values of the axial quadrupole moment, $\delta\langle\Phi|\hat{H}-\lambda\hat{Q}_{20}|\Phi\rangle=0$, and then using $q\equiv\langle\hat{Q}_{20}\rangle$. This differs from Ref. [13], where the density derivatives were determined after solving the local quasiparticle random phase approximation equations.
Since ~$\tilde{\mathcal{H}}_{1}$ up to the first order depends linearly on ~$\tilde{\mathcal{R}}_{1}$, the ATDHFB equation (5) is a linear equation for ~$\tilde{\mathcal{R}}_{1}$ in many dimensions, which necessitates an impractical inversion of an exceedingly large stability matrix [2]. To circumvent this problem, the novel method proposed here transforms the ATDHFB equation (5) into an iterative fixed-point equation,
$i\hbar\frac{\partial\mathcal{R}_{0}}{\partial{q}}=[\mathcal{H}_{0},\tilde{\mathcal{R}}_{1}^{(n)}]+[\tilde{\mathcal{H}}_{1}^{(n-1)},\mathcal{R}_{0}],$ (6)
@@
where ~$\tilde{\mathcal{H}}_{1}^{(n-1)}$ is evaluated for ~$\tilde{\mathcal{R}}_{1}^{(n-1)}$ and ~$\tilde{\mathcal{R}}_{1}^{(0)}\equiv\tilde{\mathcal{H}}_{1}^{(0)}\equiv 0$. Then, at each iteration $n=1,2,\ldots{N}$, the collective inertia can be evaluated as
@@
$\mathcal{M}^{(n)}=\frac{i\hbar}{2}\,\textrm{Tr}\left(\frac{\partial\mathcal{R}_{0}}{\partial{q}}[\mathcal{R}_{0},\tilde{\mathcal{R}}_{1}^{(n)}]\right),$ (7)
@@
where the antisymmetric matrices $Z=\varphi^{+}\tilde{\mathcal{R}}_{1}\chi$, $F=\varphi^{+}\frac{\partial\mathcal{R}_{0}}{\partial q}\chi$, and $E_{1}=\varphi^{+}\tilde{\mathcal{H}}_{1}\chi$ are defined by the standard quasiparticle ($E_{\mu}>0$) and quasihole ($E_{\mu}<0$) wave functions $\chi$ and $\varphi$, respectively, having the form
$\chi=\bigg{(}\begin{array}[]{c}A\\ B\\ \end{array}\bigg{)},\quad\varphi=\bigg{(}\begin{array}[]{c}B^{*}\\ A^{*}\\ \end{array}\bigg{)},$ (9)
with $A$ and $B$ defining the general Bogoliubov transformation. Positive quasiparticle energies ($E_{\mu},E_{\nu}>0$) are taken in Eq. (8).
@@
To proceed with the iteration in Eq. (8), $E_{1}^{(n)}$ is computed from $Z^{(n)}$, which closely resembles the standard iteration of the HFB equation. Indeed, at each iteration, we first perform the singular value decomposition (SVD) of the antisymmetric matrix $Z=U\Omega V^{+}$, where $U$ and $V$ are unitary and $\Omega$ is diagonal and non-negative; that is, $\Omega_{\mu\nu}=\delta_{\mu\nu}\omega_{\mu}$ and $\omega_{\mu}\geq 0$. Next, we observe that the eigenequation for the Hermitian matrix $\mathcal{Y}=\left(\begin{array}[]{cc}0&Z\\ Z^{+}&0\end{array}\right)$ (the first-order density matrix ~$\tilde{\mathcal{R}}_{1}$ expressed in the quasiparticle basis) takes the following form:
$\mathcal{Y}\bigg{(}\begin{array}[]{c}U\\ \pm V\end{array}\bigg{)}=\bigg{(}\begin{array}[]{c}\pm U\Omega\\ V\Omega\\ \end{array}\bigg{)}=\pm\bigg{(}\begin{array}[]{c}U\\ \pm V\\ \end{array}\bigg{)}\Omega.$ (10)
This indicates that the eigenvectors of $\mathcal{Y}$ occur in pairs of opposite eigenvalues, and the matrix $\mathcal{Y}$ can be represented as the sum of its eigenvectors [^2] Note that the eigenvectors in Eq. (10) are normalised to 2. This is so because, in the standard definition of the SVD, the columns of the unitary matrices $U$ and $V$ are normalised to 1. as follows:
$\mathcal{Y}=\tfrac{1}{2}\bigg{(}\begin{array}[]{cr}U&U\\ V&-V\\ \end{array}\bigg{)}\bigg{(}\begin{array}[]{cc}\Omega&0\\ 0&-\Omega\\ \end{array}\bigg{)}\bigg{(}\begin{array}[]{cr}U^{+}&V^{+}\\ U^{+}&-V^{+}\\ \end{array}\bigg{)}.$ (11)
Therefore, the first-order correction matrix, ~$\tilde{\mathcal{R}}_{1}=\mathcal{A}\mathcal{Y}\mathcal{A}^{+}$, for $\mathcal{A}=\left(\begin{array}[]{cc}A&B^{*}\\ B&A^{*}\end{array}\right)$, can be obtained through the inverse Bogoliubov transformation, and thus it can be represented in its final form as,
~$\tilde{\mathcal{R}}_{1}$ $=$ $\bigg{(}\begin{array}[]{cr}A_{1}&{B_{1}^{\prime}}^{*}\\ B_{1}&{A_{1}^{\prime}}^{*}\\ \end{array}\bigg{)}\bigg{(}\begin{array}[]{cc}\Omega&0\\ 0&-\Omega\\ \end{array}\bigg{)}\bigg{(}\begin{array}[]{cr}A_{1}^{+}&B_{1}^{+}\\ {B_{1}^{\prime}}^{T}&{A_{1}^{\prime}}^{T}\\ \end{array}\bigg{)}$ (18)
$=$ $\bigg{(}\begin{array}[]{cr}\rho_{1}&\kappa_{1}\\ -\kappa_{1}^{*}&-\rho_{1}^{*}\\ \end{array}\bigg{)}=\chi_{1}\Omega\chi_{1}^{+}-\varphi_{1}\Omega\varphi_{1}^{+},$ (21)
where $\rho_{1}$ and $\kappa_{1}$ are the first-order corrections to the standard density matrix and pairing tensor [2], and
$\chi_{1}$ $=$ $\begin{array}[]{l}\left(\begin{array}[]{c}A_{1}\\ B_{1}\end{array}\right)=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}AU+B^{*}V\\ BU+A^{*}V\end{array}\right),\end{array}$ (27)
$\varphi_{1}$ $=$ $\begin{array}[]{l}\left(\begin{array}[]{c}{B_{1}^{\prime}}^{*}\\ {A_{1}^{\prime}}^{*}\end{array}\right)=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}AU-B^{*}V\\ BU-A^{*}V\end{array}\right).\end{array}$ (33)
This compares perfectly well with the standard eigenstates of the static solution $\mathcal{R}_{0}$ (9), which appear in pairs with eigenvalues 0 and 1,
$\mathcal{R}_{0}$ $=$ $\bigg{(}\begin{array}[]{cr}A&B^{*}\\ B&A^{*}\\ \end{array}\bigg{)}\bigg{(}\begin{array}[]{cc}0&0\\ 0&1\\ \end{array}\bigg{)}\bigg{(}\begin{array}[]{cr}A^{+}&B^{+}\\ B^{T}&A^{T}\\ \end{array}\bigg{)}$ (40)
$=$ $\bigg{(}\begin{array}[]{cc}\rho_{0}&\kappa_{0}\\ -\kappa_{0}^{*}&1-\rho_{0}^{*}\\ \end{array}\bigg{)}=\varphi\varphi^{+}.$ (43)
We see that the zero-order density matrix $\mathcal{R}_{0}$ (40) is given as a sum of the zero-order quasihole wave functions $\varphi$ (9) with all quasioccupation factors equal to 1. Similarly, the first-order density matrix ~$\tilde{\mathcal{R}}_{1}$ (18) is given as a sum of both the first-order quasiparticle and quasihole wave functions $\chi_{1}$ and $\varphi_{1}$ (27)–(33) with the opposite quasioccupation factors of $+\omega_{\mu}$ and $-\omega_{\mu}$, respectively.
This enables an easy adaptation of the standard iteration of the HFB equation [2] to the novel iterative solution of the ATDHFB equation presented in Eq. (8). Indeed, we first determine the time-odd density matrix $\rho_{1}$ and the time-odd pairing tensor $\kappa_{1}$ in Eq. (18) based on the quasiparticle wave functions (27) and (33). Second, we employ the standard HFB algorithm [2] to determine the time-odd mean field $\Gamma_{1}$ and the time-odd pairing field $\Delta_{1}$ from $\rho_{1}$ and $\kappa_{1}$, leading to ~$\tilde{\mathcal{H}}_{1}=\left(\begin{array}[]{cc}\Gamma_{1}&\Delta_{1}\\ -\Delta_{1}^{*}&-\Gamma_{1}^{*}\\ \end{array}\right)$ and $E_{1}=\varphi^{+}\tilde{\mathcal{H}}_{1}\chi$. Third, we utilise Eq. (8) to determine the matrix $Z$ for the next iteration. And fourth, the SVD of the matrix $Z$ enables the closure of the self-consistent loop by providing the quasiparticle wave functions (27) and (33).
We see that the iterative equations (6) and (8) involve only one-body quasiparticle matrices, and the stability matrix does not need to be evaluated or inverted. Obviously, in the first iteration, in which ~$\tilde{\mathcal{H}}_{1}$ is neglected, $\mathcal{M}^{(1)}$ gives the IB collective inertia.
@@
The method was implemented in the non-relativistic density functional solver HFODD [20, 21], which solves the nuclear DFT equations by expanding the quasiparticle wavefunction on a three-dimensional Cartesian harmonic-oscillator basis and allows for the self-consistent description of nuclei with arbitrary shapes. The precision of the iterative method was verified by comparing the calculated Thouless-Valatin moments of inertia with those obtained from cranking calculations, yielding perfect agreement [(Sun25a), 18].
@@
The IB collective inertia has been widely used in multiple studies of nuclear collective motion. However, it was soon realised that the calculated values were too low to describe the data. As a result, in most applications, the IB values were multiplied by ad hoc factors of about 1.3, see, e.g., recent Refs. [22, 23, 24] and references cited therein. Those were purely empirical factors meant to compensate for the unknown effects of the time-odd mean fields. In the present work, having a rapid and efficient method to include those effects, we microscopically determined the ratios of the ATDHFB and IB moments of inertia and compared the full ATDHFB results with the data.
To this end, we performed systematic ATDHFB calculations for the Gd, Dy, Er, Yb, Hf, W, and Os isotopes with neutron numbers between 82 and 126. Two different Skyrme functionals, SkM* [25] and UNEDF1 [26], and one Gogny functional D1S [27] were used for this purpose. For SkM*, the volume pairing with the neutron and proton strengths of --178.83 and --211.20 MeV fm3 [12] were used, respectively. For UNEDF1, the mixed pairing of the type $V_{0}^{t}(1-\rho/\rho_{\textrm{sat}})$ was used with the neutron and proton parameters $V_{0}^{n}=-223.278$ and $V_{0}^{p}=-247.896$ MeV fm3, and with $\rho_{\textrm{sat}}=0.32~{}\textrm{fm}^{-3}$. Since the time-odd sectors of those Skyrme functionals were originally not adjusted to data, here we fixed them by employing the adjustments performed in Ref. [28].
@@
In Fig. 2, the distribution of the residuals between the calculated moments of inertia and the experimental values $(\mathcal{I}^{\textrm{Th.}}-\mathcal{I}^{\textrm{Expt.}})$ of all the experimentally available isotopes between $N$ = 86 and 122 are presented for SkM*, UNEDF1, and D1S functionals, respectively. The shift between the distribution of the residuals for IB and ATDHFB moments of inertia can be clearly observed, which shows the importance of the time-odd mean field.
Sudden increase of the deformation in the Hf, W and Os isotopes around $N=100$, predicted by UNEDF1 and D1S, is not visible in the experimental data, possibly because triaxiality is neglected in our calculations. Therefore, in the following we present the statistical analysis corresponding to both excluding and including these nuclei, with the latter results given in parentheses.
For SkM*, the mean value of the residual between the IB moment of inertia and the experimental data is --2.55 (--2.21) 2superscriptPlanck-constant-over-$\hbar^{2}$/MeV, that is, for this functional, the IB formula underestimates the moment of inertia. On the contrary, an overestimation is predicted by the ATDHFB method with a mean value of 2.54 (3.06) 2superscriptPlanck-constant-over-$\hbar^{2}$/MeV. In both cases, the distributions are fairly narrow with standard deviations of 4.33 (4.21) 2superscriptPlanck-constant-over-$\hbar^{2}$/MeV and 3.36 (3.48) 2superscriptPlanck-constant-over-$\hbar^{2}$/MeV for IB and ATDHFB, respectively. For D1S, the IB formula also underestimates the moment of inertia while the ATDHFB method overestimates them, but the distributions of the residual are wider than for SkM*. For UNEDF1, wider distributions are also found but both the IB formula and ATDHFB method overestimate the experimental values.